2. King Supply makes four different types of plumbing fixtures: W, X, Y and Z. The contribution margins for these products are: $70 for Product W, 560 for Product X, $90 for Product Y and S100 for Product Z. Fixed overhead is estimated at S5,500 per week. The manufacture of each fixture requires four machines, Machines #1, 2, 3 and 4, Each of the machines is available for 40 hours a week and there is no setup time required when shifting from the production of one product to any other. The processing requirements to make one unit of each product are shown in the table. Weekly product demand for the next planning period has been forecasted as follows: 70 WS, 60 50 Ys and 30孫 In all the questions below show calculations as needed Note: 1. Fixtures W, X,Y and Z are the Products of King Supply company (equivalent to products Contribution Margins of the products are already given C. D in the homework problem 2. Processing Time (Minutes Per Fixture Fixture MachineMachine 2 Machine 3 Machine 4 20 15 A) Use the information in above table, which machine is the bottleneck operation? (5 Points) A) Machine 1 B) Machine 2 C) Machine 3 D) Machine 4 Answer: B) Use the information in the above table and using the bottleneck method, in what sequence should products be scheduled for production? (5 Points) A)Z, Y, X, W B) X, W, Z, Y C) Z, Y, W, X D) X, Y, Z, W Answer:
Answer #1

A) Answer: A) Machine 1

Workload of machine 1 = 10*70+5*60+20*50+15*30 = 2450 minutes

Workload of machine 2 = 15*70+10*60+5*50+5*30 = 2050 minutes

Workload of machine 3 = 10*70+15*60+5*50+5*30 = 2000 minutes

Workload of machine 4 = 5*70+10*60+10*50+10*30 = 1750 minutes

Available capacity per machine = 40 hours* 60 minutes per hour = 2400 minutes

Workload of machine 1 is more than its capacity. Therefore, Machine 1 is the bottleneck.

B) Answer: B) X,W, Z, Y

Using bottleneck method, we need to determine the profit contribution per minute of bottleneck resource (Machine 1)

Profit contribution per minute of bottleneck resource for product W = 70/10 = $ 7

Profit contribution per minute of bottleneck resource for product X = 60/5 = $ 12

Profit contribution per minute of bottleneck resource for product Y = 90/20 = $ 4.5

Profit contribution per minute of bottleneck resource for product Z = 100/15 = $ 6.7

The products are scheduled in order of their profit contribution per minute of bottleneck resource. So the optimal schedule is: X,W, Z, Y

C) Answer: C) 70 W, 60 X, 47 Y, 30 Z

Product X, W and Z can be produced upto their respective demand level. Remaining time on Machine 1 = 2400 - (5*60+10*70+15*30) = 950 minutes

Quantity of Product Y that can be produced in remaining time = 950/20 = 47.5 ~ 47

D) Answer: B) greater than $ 10000 but less than or equal to 11000

Profit = Profit contribution from products - Fixed overhead cost

= (70*70+60*60+47*90+30*100)-5500

= $ 10,230